Circular No. 6309

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 T. Augusteijn and G. van de Steene, European Southern Observatory;

D. A. Frail, Very Large Array, National Radio Astronomy Observatory; J. van Paradijs, University of Alabama at Huntsville and University of Amsterdam; C. Kouveliotou, Universities Space Research Association; G. J. Fishman, Marshall Space Flight Center, NASA; and W. H. G. Lewin, Department of Physics, Massachusetts Institute of Technology, report: “By comparing the radio position with the digitized Sky Survey, we find a faint source that falls within ∼ 2′′ of the position of the radio source in the error box of GRO J1744–28 (cf. IAUC 6307). From an image taken on Jan. 21.360 UT, we find V ∼ 19.8. From an image taken on Feb. 9.322, we find R = 18.1 ± 0.1. If GRO J1744–28 is a normal, low-mass, x-ray binary, one would expect its absolute visual magnitude to be M_{V} ∼ −3. If the radio source is the counterpart of the x-ray source, at a distance of 7 kpc, the visual extinction of the source is at least 8 magnitudes, corresponding to a column density N_{H} > 2 × 10^{22} cm^{−2}. However, given the optical magnitudes of the object and the fact that the source can be seen on the digitized Sky Survey, this indicates that, if the source is a low-mass xray binary, it is not very strongly reddened. This is also supported by comparison to other stars in the field seen in the V image and Gunn z images taken shortly afterwards. If N_{H} < ∼ 10^{22} cm^{−2} (cf. IAUC 6291), the weakness of the optical counterpart may require beaming of the x-ray emission.”

T. Strohmayer, Universities Space Research Association; and K. Jahoda and F. Marshall, Goddard Space Flight Center, NASA, communicate for the XTE science and instrument teams: “Additional proportional counter array (PCA) scanning observations of GRO J1744–28, obtained since the first observations by XTE on Jan. 18–19 (cf. IAUC 6291), have been analyzed the further to constrain the object’s position. The resulting position obtained by averaging the results of seven independent determinations is α = 17^{h}44^{m}34^{s}.3 ± 2^{s}.8, δ = −28^{o}45′22′′ ± 47′′ (equinox 2000.0). The quoted uncertainties are the standard deviations among the positions determined from seven independent scanning observations. These uncertainties are consistent with our current understanding of the possible systematic errors.”

1996 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green